OpenStudy (anonymous):

A bank gives 3% interest on its savings accounts, compounded monthly. If one opens a savings account and deposits $ 300 each month for a total of 11 years, how much money will be in the account after the last deposit is made?

OpenStudy (amistre64):

3% is spread over 12 months; and the interest as calculated and added each month

OpenStudy (amistre64):

B[n] = B_[n-1] + B{n-1}(.03/12) B[n] = B_[n-1] (1+(.03/12)) B[n] = B[0](1+(.03/12))^n where n is measured in months

OpenStudy (amistre64):

gotta add in the 300 per month tho ... doh!!

OpenStudy (amistre64):

B[n] = B_[n-1] + B{n-1}(.03/12) B[n] = B_[n-1] (1+(.03/12)) + 300n hmmm....

OpenStudy (amistre64):

B[n] = B_[n-1] (1+(.03/12)) + 300n B[n-1] = B_[n-2] (1+(.03/12)) + 300(n-1) B[n] = (B_[n-2] (1+(.03/12)) + 300(n-1)) (1+(.03/12)) + 300n B[n] = B_[n-2] (401/400))^2 + (401/400)300n- (401/400)300 + 300n B[n] = B_[n-2] (401/400))^2 + 300(401/400)n- (401/400) + n) \[B_n = B_{n-2} (401/400)^2 + 300(\frac{401}{400}n-\frac{401}{400} + n)\] oy, this would work out better on paper, and also if i knew an easier trick :)